# American Sign Language Translation ASLP American Sign Language Production

### v2

@@
Sen Fang1, Lei Wang2,3, Ce Zheng4, Chunyu Sui5, Mingyu Zhao1, Yapeng Tian6, Chen Chen7
1 Rutgers University, 2 Australian National University, 3 Data61/CSIRO, 4 Carnegie Mellon University
5 Columbia University, 6 The University of Texas at Dallas, 7 University of Central Florida
@@ Abstract
In this paper, we propose SignLLM, a multilingual Sign Language Production (SLP) large language model, which includes two novel multilingual SLP modes MLSF and Prompt2LangGloss that allow sign language gestures generation from query texts input and question-style prompts input respectively. Both modes can use a new RL loss based on reinforcement learning and a new RL module named Priority Learning Channel. These RL components can accelerate the training by enhancing the model’s capability to sample high-quality data. For SignLLM’s training, we introduce Prompt2Sign, a comprehensive multilingual sign language dataset, which builds from public data, including American Sign Language (ASL) and seven others. This dataset standardizes information by extracting pose information from sign language videos into a unified compressed format. We extensively evaluate SignLLM, demonstrating that our model achieves state-of-the-art performance on SLP tasks across eight sign languages.
Figure 1: Overview: (Left) Major components (e.g., Refer to caption, Refer to caption, Refer to caption, etc.) of our Prompt2Sign dataset. Compressed Pose is reprocessed pose data that is suitable for training, we use public sign language videos to produce compressed pose data in our predefined format; (Right) Our proposed SignLLM aims to generate sign language poses for various application scenarios [12, 6, 110, 105].
@@ 1 Introduction
To overcome these challenges, we introduce SignLLM, a large multilingual Sign Language Production (SLP) model developed based on our Prompt2Sign dataset. It can produce the sign language representation of eight languages from texts or prompts. Our SignLLM has two distinct modes: (i) Multi-Language Switching Framework (MLSF), which allows multiple sign languages production in parallel by dynamically adding encoder-decoder groups. (ii) Prompt2LangGloss, allowing SignLLM to support static single-set encoder-decoder generation. Fig. 1 (Right) shows our model inputs and outputs, “thank you” is the input of mode (i), and “how to sign ‘thank you’ in ASL?” is the input of mode (ii). Two multilingual SLP modes deal with different use cases: The MLSF is an efficient mode without semantic confusion, like a dictionary/drawer; The Prompt2LangGloss is a user-friendly mode, like a LLM, it aims to understand complex natural language input. They are like two evolutionary branches diverged from the same species. To address the problem of extended training time caused by more languages and a larger dataset, we utilize the Reinforcement Learning (RL) Loss concepts to quantify the quality of each training batch and prioritize valuable batches through the Priority Learning Channel module.
@@
  • A comprehensive multilingual sign language dataset, named Prompt2Sign, featuring an expanded vocabulary and covering eight languages, is introduced. This dataset is designed for compatibility with a wide range of seq2seq models.
  • A large multilingual Sign Language Production (SLP) model with two distinct modes—MLSF for handling text query inputs and Prompt2LangGloss for processing question-style prompts—is introduced. Our method, SignLLM, achieves state-of-the-art performance across eight sign languages in SLP tasks.
  • We introduce a novel reinforcement learning-based loss function, along with a functional module named PLC, as a training strategy for sign language models, designed to reduce training time and computational costs.
@@ 3 Our Benchmark: Prompt2Sign
Dataset Range. We have cited the sources of our publicly available data, and some of the more popular works were not considered due to their limited accessibility and potential usage restrictions. Additionally, while there are other multilingual datasets available [29, 101, 55, 33], they may not possess the same level of comprehensiveness as ours. For example, [29] and [101] translate two types of sign language videos into spoken language (SLT), while our work is from spoken language to videos (SLP). We aim to establish a robust multilingual SLP method with a dataset that supports a broader range of application scenarios. For example, [29] is limited to Bible translation, and [101] focuses solely on cross-lingual SLT. In contrast, our work covers a broader range of scenarios and comprehensively addresses the SLP task. Additionally, some datasets, such as [55] and [33], are restricted to multilingual dictionary forms.
@@ 4.4 Reinforcement Learning Training Strategy
Table 2: American Sign Language Production (ASLP): Comparison of SignLLM variants with baseline on Text to Pose task by using our Prompt2Sign ASL part. The improvement (+num) is relative to the latest work [22].
@@ 5 Experiments and Discussions
Metrics. We mainly use 3 metrics: (i) BLEU-n score measures the similarity between machine-generated translations and reference translations based on n-grams, the closer the predicted result is to the input (reference), the higher the value. BLEU-n [60] means that n words are used as the basic computing unit, and the higher the n, the higher the fluency requirement. (ii) ROUGE score [49] is similar to BLEU, but it is more concerned with consistency and coverage. It indicates better agreement between the generated and reference texts, indicating a more accurate and comprehensive summary. (iii) DTW score underpinned by dynamic programming principles [3], is employed to ascertain the smallest manipulation distance between clips and sentences; the lower, the better. More model information and experiments can be found in supplementary materials.
@@
Base 10.96 14.68 22.49 40.29 44.13 9.27 10.72 18.64 39.88 40.39
Prompt2LangGloss 15.78 22.66 30.27 48.30 50.88 16.32 20.27 28.70 47.89 48.18
RL Loss 18.33 24.28 32.18 48.83 52.61 13.26 17.33 24.51 40.76 40.95
RL Loss & PLC 18.77 25.42 32.44 52.66 61.44 14.13 19.08 25.72 43.40 51.57
Table 5: Comparison of Different Modules: SignLLM-40M-Base results for Text to Pose task on the ASL part of Prompt2Sign, with multiple data augmentation techniques. Base: MLSF with Normal MSE Loss. Prompt2LangGloss: Prompt2LangGloss with normal MSE Loss. PLC: Priority Learning Channel. Top performances are highlighted in bold, while second top performances are underlined.
@@ 5.1 Quantitative Evaluation
Back Translation. Back translation means translating generated sign language videos back into spoken language sentences. These sentences are compared with the original input sentences to evaluate the translation quality. The task is widely adopted to evaluate Sign Language Production (SLP) as it can indicate the accuracy of the produced sign language videos [71], translation models are trained on the corresponding language skeletal pose data.
@@ 5.2 Ablation Evaluation
Figure 4: RL Training Efficiency Analysis: Comparison of different Settings on DTW values (the lower the better) at different epoch times. Left Y-axis: Value of DTW. Right Y-axis: Value of Loss. Prompt: Prompt2LangGloss mode.
Performance. Table 5 indicates our last two innovative strategies (RL Loss and PLC module) significantly improve the model’s performance, and both strategies contribute to substantial improvements. We find that with sufficient data, LangGloss can outperform the base model without enhancements, showing good usability. So, we conducted more research on this in the supplementary materials. In addition, despite the demonstrable improvement, RL Loss facilitates in the development set (showing improvements of 7.37 in BLEU-4 and 8.48 in ROUGE scores), its effectiveness is less apparent in the test set. However, a completely different picture emerges when we combine RL Loss with the PLC. This innovative integration improves the ROUGE score by 25.9% on the test set (from 40.95 to 51.57), showcasing the positive impact of our proposed technique. These results show that PLC has successfully enhanced the adaptability of the original RL Loss to unknown environments.

### v3

@@
Sen Fang1, Chen Chen2, Lei Wang3,4, Ce Zheng5, Chunyu Sui6, Yapeng Tian7
1 Rutgers University, 2 University of Central Florida, 3 Australian National University, 4 Data61/CSIRO
5 Carnegie Mellon University, 6 Columbia University, 7 The University of Texas at Dallas
@@ Abstract
In this paper, we propose SignLLM, a multilingual Sign Language Production (SLP) large language model, which includes two novel multilingual SLP modes MLSF and Prompt2LangGloss that allow sign language gestures generation from query texts input and question-style prompts input respectively. Both modes can use a new RL loss based on reinforcement learning and a new RL module named Priority Learning Channel. These RL components can accelerate the training by enhancing the model’s capability to sample high-quality data. To train SignLLM, we introduce Prompt2Sign, a comprehensive multilingual sign language dataset, which builds from public data, including American Sign Language (ASL) and seven others. This dataset standardizes information by extracting pose information from sign language videos into a unified compressed format. We extensively evaluate SignLLM, demonstrating that our model achieves state-of-the-art performance on SLP tasks across eight sign languages.
Figure 1: Overview: (Left) Major components (e.g., Refer to caption, Refer to caption, Refer to caption, etc.) of our Prompt2Sign dataset. Compressed Pose is reprocessed pose data that is suitable for training, we use public sign language videos to produce compressed pose data in our predefined format; (Right) Our proposed SignLLM aims to generate sign language poses for digital human or avatar generation [12, 6, 111, 106].
@@ 1 Introduction
To overcome these challenges, we introduce SignLLM, a large multilingual Sign Language Production (SLP) model developed based on our Prompt2Sign dataset. It can produce the sign language representation of eight languages from texts or prompts. Our SignLLM has two distinct modes: (i) Multi-Language Switching Framework (MLSF), which allows multiple sign languages production in parallel by dynamically adding encoder-decoder groups. (ii) Prompt2LangGloss, allowing SignLLM to support static single-set encoder-decoder generation. Fig. 1 (Right) shows our model inputs and outputs, “thank you” is the input of mode (i), and “how to sign ‘thank you’ in ASL?” is the input of mode (ii). Two multilingual SLP modes deal with different use cases: The Multi-Language Switching Framework (MLSF) is an efficient mode without semantic confusion, like a dictionary/drawer; The Prompt2LangGloss is a user-friendly mode, like a LLM, it aims to understand complex natural language input. To address the problem of extended training time caused by more languages and a larger dataset, we utilize the Reinforcement Learning (RL) Loss concepts to quantify the quality of each training batch and prioritize valuable batches through the Priority Learning Channel.
@@
  • A comprehensive multilingual sign language dataset, named Prompt2Sign, featuring an expanded vocabulary and covering eight languages, is introduced. It is designed for broader seq2seq compatibility.
  • A large multilingual Sign Language Production (SLP) model with two distinct modes—MLSF for handling text query inputs and Prompt2LangGloss for processing question-style prompts—is proposed. Our method, SignLLM, achieves state-of-the-art performance across eight sign languages in SLP tasks.
  • We present a novel reinforcement learning-based loss function, along with a functional module named Priority Learning Channel (PLC), as a training strategy for sign language models, designed to reduce training time and computational costs.
@@ 3 Our Benchmark: Prompt2Sign
Dataset Range. We choose How2Sign [19], PHOENIX-14T [26], KSL-guide [32], Signsuisse [59] (contains 3 languages), LSA64 [69], AuTSL [81], and some of the more popular works were not considered due to their limited accessibility and potential usage restrictions. Additionally, some available multilingual datasets [30, 102, 56, 34] may not possess the same level of comprehensiveness as ours. For example, [30] and [102] translate two types of sign language videos into spoken language (SLT), while our work is from spoken language to videos (SLP). We aim to establish a robust multilingual SLP method with a dataset that supports a broader range of application scenarios. For example, [30] is limited to Bible translation, and [102] focuses solely on cross-lingual SLT. In contrast, our work covers a broader range of scenarios and comprehensively addresses the SLP task. Additionally, some datasets, such as [56] and [34], are restricted to multilingual dictionary forms.
@@ 4.4 Reinforcement Learning Training Strategy
Table 2: American Sign Language Production (ASLP): Comparison of SignLLM variants with baseline on Text to Pose task by using our Prompt2Sign ASL part. “-” : The NSA [2] and NSLP-G [39] have not been tested on the dev set, and there is no source code. The improvement (+num) is relative to the latest work [22].
@@ 5 Experiments and Discussions
Metrics. (i) BLEU-n score measures the similarity between machine-generated translations and reference translations based on n-grams, the closer the predicted result is to the input (reference), the higher the value. BLEU-n [61] means that n words are used as the basic computing unit, and the higher the n, the higher the fluency requirement. (ii) ROUGE score [50] is similar to BLEU, but it is more concerned with consistency and coverage. It indicates better agreement between the generated and reference texts, indicating a more accurate and comprehensive summary. (iii) DTW score underpinned by dynamic programming principles [3], is employed to ascertain the smallest manipulation distance between clips and sentences; the lower, the better.
@@
Base + Normal MSE Loss 10.96 14.68 22.49 40.29 44.13 9.27 10.72 18.64 39.88 40.39
Base + RL Loss 18.33 24.28 32.18 48.83 52.61 13.26 17.33 24.51 40.76 40.95
Base + RL Loss & PLC 18.77 25.42 32.44 52.66 61.44 14.13 19.08 25.72 43.40 51.57
Base + MSE + Prompt2LangGloss 15.78 22.66 30.27 48.30 50.88 16.32 20.27 28.70 47.89 48.18
Base + MSE + MLSF 16.44 23.79 32.32 50.97 53.44 17.17 21.25 30.12 50.25 50.47
Table 5: Comparison of Different Modules: Base: SignLLM-40M-Base results for Text to Pose task on the ASL part of Prompt2Sign. PLC: Priority Learning Channel. Top performances are highlighted in bold, while second top performances are underlined.
@@ 5.1 Quantitative Evaluation
Back Translation. Back translation means translating generated sign language videos back into spoken language sentences. These sentences are compared with the original input sentences to evaluate the translation quality. The task is widely adopted to evaluate Sign Language Production (SLP) as it can indicate the accuracy of the produced sign language videos [72], our translation models [15] are trained on the corresponding language Prompt2Sign pose data, the performance is between bleu-4 22.3 and 27.6, it is comparable to the translation model used in previous work.
@@ 5.2 Ablation Evaluation
Figure 4: RL Training Efficiency Analysis: Comparison of different Settings on DTW values (the lower the better) at different peroid (every 30 epochs a peroid). Left Y-axis: Value of DTW. Right Y-axis: Value of Loss. Prompt: Prompt2LangGloss mode.
Performance. Ablation results in Table 5 indicates our four innovative strategies (Prompt2LangGloss, MLSF modes, RL Loss and PLC module) significantly improve the model’s performance, and four strategies contribute to substantial improvements. When replacing the standard MSE Loss with RL Loss, we observe a significant improvement in the model’s performance. These results show that PLC has successfully enhanced the adaptability of the original RL Loss to unknown environments. Incorporating the PLC module further enhances these gains. We also note that the base model with MSE, when upgraded with Prompt2LangGloss, achieves higher scores with the new mode, slightly underperforming compared to the MLSF mode but far surpassing the baseline model in the first row. This indicates that, with sufficient data, the Prompt2LangGloss mode demonstrates good usability, and we explore this further in the supplementary materials.